# SmallPlan: Leverage Small Language Models for Sequential Path Planning with Simulation-Powered, LLM-Guided Distillation

### v3

@@ 5.2 Results
Framework Model Fine-tune SR % $\uparrow$ SPL % $\uparrow$ Dist. $\downarrow$ Trials $\downarrow$ Acts. Strategy (High/Objs) $\downarrow$ Pretrained / MoMaLLM GPT-4o - 97.14 65.92 18.01 0.39 7.30/3.49 Pretrained Qw2.5-3B - 0.00 0.00 - - - SmallPlan w/o RL Qw2.5-3B SFT only 93.14 58.15 21.42 1.60 9.74/4.23 SmallPlan Qw2.5-3B RL-SFT $+$ 78.86 59.31 11.49 1.05 6.01/2.69 SmallPlan Qw2.5-3B RL-SFT * 83.43 48.07 19.87 1.04 8.93/2.48 Pretrained Phi4-mini - 70.86 41.57 17.84 3.39 10.68/2.18 SmallPlan w/o RL Phi4-mini SFT only 91.43 58.40 18.26 0.88 7.94/3.17 SmallPlan Phi4-mini RL-SFT $+$ 85.14 60.33 13.52 0.91 6.41/2.66 SmallPlan Phi4-mini RL-SFT* 94.29 64.00 16.72 0.49 6.55/2.85
Table 2: Break-down SR % and SPL % for each test scenes. Numbers rounded to one decimal digit. $+$ and * are similar to Table 1
Model Beechwood 0 Benevolence 1 Benevolence 2 Merom 1 Pomaria 1 Pomaria 2 Wainscott 0 SR SPL SR SPL SR SPL SR SPL SR SPL SR SPL SR SPL GPT-4o 100 56.8 100 96.6 96 53.9 96 62.9 100 67.4 100 71.9 88 51.9 Qw2.5-3B 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 Qw2.5-3B SFT 100 53.6 100 90.2 92 47.5 88 47.4 100 65.9 100 58.9 72 43.5 Qw2.5-3B RL-SFT+ 76 45.4 96 90.0 68 46.7 68 49.9 92 70.8 92 67.3 60 45.2 Qw2.5-3B RL-SFT 96 52.0 100 81.1 60 24.1 76 33.1 92 50.8 92 63.3 68 32.1 Phi4-mini 68 31.3 100 80.4 64 28.0 60 24.0 72 44.4 80 50.3 52 32.6 Phi4-mini SFT 92 50.9 100 90.7 88 46.5 84 47.1 92 63.8 96 61.5 88 48.4 Phi4-mini RL-SFT + 92 57.2 96 87.9 76 45.6 80 46.3 92 72.8 92 71.6 68 40.9 Phi4-mini RL-SFT 100 60.5 100 91.5 88 53.6 92 53.4 96 68.2 100 73.2 84 47.6
SmallPlan for Sequential Path Planning.
While GPT-4o consistently achieves the best results, our findings in Tables 1 and 2 show that significantly smaller models can also perform competitively on sequential path planning tasks. Both fine-tuning strategies, SFT-only and RL-interleaving-SFT, enhance the SLMs’ performance. We observe that the pretrained Qwen-2.5-3B could not follow the required response format provided in our prompt, and sometimes output Chinese answers instead of English. However, applying SFT helps reduce this issue. The experiment also shows that the fine-tuned SLMs perform well when the start poses, target objects, and object distributions modified after each runs. Additionally, we suggest that the better outcome of Phi-4-mini comes from the fact that its pretrained version is designed for complex reasoning [54].
Side Factor Control via RL.
@@
In Tables 1 and 2, we show that the SLMs adapt to different reward penalties. For RL-SFT +, we set the $\lambda_{\text{efficiency}}$ twice to that of RL-SFT (0.6 vs. 0.3). There exist trade-off between SR and Dist., as we observe that the SLMs tend to prematurely terminate tasks by calling done() instead of continuing to explore the environment. However, our approach paves the way for adapting to different constraints and settings. While SFT effectively transfers reasoning capabilities from LLMs to SLMs, it does not account for critical deployment constraints such as energy consumption. We observe that while SFT is capable of re-prompting until the maximum number of trials is reached, RL allows the model to terminate the process early, which can be seen as Act.. In real-time robotic applications on edge devices, each prompt to an LLM incurs significant energy cost and action execution overhead, making efficient operation a key challenge. RL plays a crucial role in controlling side factors, enables more resource-efficient, energy-aware planning for deployment.
Table 3: Evaluation of SmallPlan on unseen layouts with balanced reward.
Model Fine-tune SR % $\uparrow$ SPL % $\uparrow$ Dist. $\downarrow$ Trials $\downarrow$ Strategy Qw2.5-3B SFT-only 60.85 47.56 11.23 5.07 Qw2.5-3B RL-SFT 76.57 55.42 14.42 2.08 Phi4-mini SFT-only 89.71 63.92 14.15 1.07 Phi4-mini RL-SFT 92.29 64.10 15.72 0.49
Unseen Evaluation. In our original setup, although object distributions and target goals are randomized, many training and test scenes share similar environment layouts, which can limit the evaluation of a model’s true generalization capability. To better isolate this effect, we construct a new training set by removing 4 scenes that have the same layout with any of the 7 test scenes. We observe that RL-SFT outperforms SFT across most evaluation metrics. While SFT performs better when layout correlations exist (as in the 8-scene training setup), RL-SFT demonstrates stronger generalization to out-of-distribution scenes. This suggests that SFT tends to rely more on memorization, whereas RL enables the model to learn more adaptable and transferable behaviors.
@@
Response Format Correction with SFT beforehand. In Section 4.2, we mentioned that SFT is applied before the main fine-tuning process. Specifically, we first apply LoRA and 4-bit quantization to the candidate SLM, then SFT with a small sample of response output from the teacher LLM. This SFT stage fits with the LoRA setup as it helps the SLMs to highlight the important parameters for the target task. In our experiment, removing this step results in unstable training dynamics and incoherent responses, as the scalar reward from RL alone is insufficient to guide learning. Thus, we observe that fine-tuning on top of this SFT pre-adapted model is significantly more effective than pretrained one, as the SFT stage provides a strong prior for LoRA adaptation.
Table 4: Success Rate (%) of the pre-trained and fine-tuned SLMs across 4 different logic text-based games. Model notations are similar to Table 1.
Model Fine-tune Strategy Island Password StringSearch OrderingText Mean SR % $\uparrow$ Easy Med Easy Med Easy Med Easy Med Easy Med Qw2.5-3B - 3.3 0.1 39.8 11.5 11.9 3.5 21.3 2.5 19.1 4.4 Qw2.5-3B SFT-only 1.0 0.0 43.6 14.4 14.7 4.5 14.5 1.5 18.5 5.1 Qw2.5-3B RL-SFT 6.6 0.1 42.1 14.2 8.0 5.7 11.4 2.4 17.0 5.6 Phi4-mini - 7.6 0.6 46.2 20.5 22.1 6.0 33.0 4.6 27.2 7.9 Phi4-mini SFT-only 10.4 0.9 46.6 20.7 21.6 6.7 27.5 4.7 26.5 8.3 Phi4-mini RL-SFT 10.7 0.9 46.3 20.6 20.4 8.3 33.8 4.8 27.8 8.7
@@
Out-of-domain Analysis. Table 4 evaluates the logical reasoning capabilities of the SLMs on out-of-domain tasks beyond path planning. These tasks involve different types of sequential decision-making, allowing us to assess general reasoning ability. Results demonstrate that our fine-tuned SLMs perform on par with, and sometimes exceed, their pretrained versions. This suggests that distillation from a superior LLM (GPT-4o) can potentially improve the general reasoning ability of smaller models across diverse domains. Given the complexity of cross-domain adaptation, we leave further investigation to future research works. This experiment merely highlights that our fine-tuned SLMs maintain strong generalization and are not overfitting to path planning tasks.
6 Conclusion
We introduce SmallPlan, a novel LLM-guided distillation framework that leverages SFT and RL to fine-tune SLMs for high-level path planning in complex environments. In our experiments, we discuss some important findings on training strategies and out-of-domain analysis. In Section 7 below, we discuss limitations and challenges that we encountered, and provide some ideas for future works. We show that the fine-tuned SLMs reaches competitive performance compared to superior LLM such as GPT-4o. SmallPlan paves the way for practical deployment of language-driven planning systems on edge devices in real-world robotics.
7 Limitations
This section outlines several limitations and open challenges that we encountered during framework design and experiment.
Model Selection and DeepSeek.
Alongside Qwen-2.5-3B and Phi4-mini, we tried DeepSeek-R1-Distill-Qwen-1.5B as one candidate SLM for SmallPlan. However, this version encounters same issue with Qwen-2.5-3B mentioned in Section 5, outputting Chinese answers instead of English with poorer outcomes compared to the Qwen-2.5-3B. Aside having fewer params compared to its larger counterparts, we suggest that there might be language imbalance in the training dataset for Qwen and DeepSeek, which could be resolved in the future.
SFT and Distillation with Ground Truth Labels.
@@ 7 Limitations
Standard distillation with SFT typically combines the teacher model’s outputs with supervision from true labels, enabling the student model to both inherit the teacher’s knowledge and align with task objectives. In our setting for SFT, however, our Cross Entropy Loss relies solely on the teacher LLM’s outputs. This is because our task - path planning in a simulation environment - provides scalar reward signals instead of differentiable ground-truth labels. These rewards cannot be directly used in standard loss functions for backpropagation. In our approach, this issue is resolved by leveraging not only SFT but RL, which takes simulation feedback alongside the teacher LLM’s outputs to optimize the candidate SLM. Future work may explore resolutions to further enhance the SFT process.
Choice of RL Algorithm.
@@
Recent works introduce different alternatives to PPO. For instance, Direct Preference Optimization (DPO) [55] is an outstanding approach that eliminates the need for an explicit teacher model, making the training process more efficient while often achieving performance comparable to PPO. However, DPO offers much less freedom in customizing the reward function to control other important factors. The DPO’s objective is fixed by pair-wise preferences and a KL-regularization term, makes it difficult to integrate the additional reward components described in Section 4.2. In contrast, PPO accepts any scalar reward; we can mix multiple objectives, tune their weights, and monitor their individual effects during training. These reasons make PPO more feasible for our approach. However, future efforts may revisit DPO or other newer algorithms in search of a better RL strategy.

### v5

@@ IV-B3 SmallPlan Training Strategy
[…]
6:end for
7:Stage 2: Interleaving RL–SFT Training
8:for $epoch\leftarrow 1$ to num_epochsnum\_epochs do
9:  for task_idtask\_id in taskstasks do
10:   $info\leftarrow\textsc{GetEnvInfo}(task\_id)$
11:   $prompt\leftarrow\textsc{AddTo}(prompt\_template,info)$
[…]
14:   $todo\_actions\leftarrow\textsc{Parse}(slm\_ans)$
15:   $sim\_feedback\leftarrow\textsc{Execute}(todo\_actions)$
[…]
18:  end for
19:end for
In the first stage, the candidate SLM is supervised fine-tuned with a small set of response samples from the teacher LLM. Pretrained SLMs often fail to produce outputs in a consistent format, even with well-crafted instruction prompts, making it difficult to parse their responses into executable commands. This step thus helps the SLM learns the correct output format from the LLMs. This few-shot SFT strategy also fits with the LoRA setup as it helps the SLMs to highlight the important parameters for the target task. To speed up this stage, we prepare these sample subsets offline beforehand and train SFT with these samples.
In the second stage, we train the SLM using an interleaved combination of RL and SFT. At runtime, we collect environment data including detected objects, their distances, and their relations from the scene graph and Voronoi graph, then parse this information into the prompt template to build a complete prompt as input for LLM and SLM to response. Here, SFT acts as a distillation mechanism for transferring reasoning behavior from the LLM, while RL - using our $r_{\text{action}}$ reward - adjusts the model’s planning behavior and controls task-related factors. This stage can also be configured to use only SFT, as distillation alone can sometimes be enough to achieve strong performance. In our experiments in Section V, we compare different training strategies to evaluate their effectiveness. Our framework, however, strongly discourages removing the SFT and relying solely on RL, as this would prevent the SLM from effectively learning from the LLM.
V Experiments
We study the following key points: (1) The effectiveness of SmallPlan in sequential path planning tasks; (2) The impact of different training strategy and trade-offs when using RL to control side factors; (3) the role of SFT during pre-training in enhancing subsequent RL training; and (4) How the trained SLMs generalize to out-of-distribution tasks.
V-A Experiment Setup
V-A1 Model Selection
We investigate various SLMs, including Qwen-2.5 (3.1B params) [47], Llama-3 (3.2B) [48], and Phi4-mini (3.8B) [49]. We used the dynamic 4-bit quantized versions provided by Unsloth AI. These models consume approximately 4GB of memory for inference. For smaller SLMs around 1B params, we discuss their performances in Section VI. We compare SmallPlan’s performance with the state-of-the-art MoMaLLM [14], which utilized GPT-4 as the language model for task planning.
V-A2 Simulation Setup and Dataset
We conduct experiments using the iGibson simulation platform [50, 51] with the Fetch robot, adopting the setup from [14]. The environment is configured based on the iGibson Challenge data split with 8 train scenes for development and 7 test scenes for evaluation. Each run within a scene randomizes the start poses, target objects, and object distributions, resulting in a diverse spectrum of task scenarios and action sequences. We evaluate our method over 25 run trials per test scene, yielding a total of 175 challenging and previously unseen task executions.
V-A3 Out-of-domain Reasoning Analysis Setup
We conduct an out-of-domain experiment to assess the risk of overfitting in our fine-tuned SLMs. We use a text-based interactive game benchmark [52], designed to evaluate various logical reasoning skills of language models. Specifically, we compare the trained SLMs with their pretrained versions on 4 different games introduced in this benchmark.
V-A4 Metrics.
We evaluate SmallPlan using 4 metrics:
Task Success Rate (SR %): The percentage of successfully completed tasks out of all attempted trials. For out-of-domain analysis, this corresponds to the percentage of solved challenges out of all attempts.
Success weighted by Path Length (SPL %): The ratio of the shortest possible path length to the actual path taken, weighted by task success.
Distance Traveled (Dist.): The total path length covered by the agent across all task executions.
Average Retrials (Retrials): The average number of retries attempted by the agent per run per scene.
V-B Results
Figure 3: The mean SR (%) of the pre-trained and fine-tuned SLMs for the text-based games on Easy and Medium modes. There are 4 different games, with each game played 1,000 times with different sub-challenges provided by [52].
V-B1 SmallPlan for Sequential Path Planning
While GPT-4o consistently achieves the highest SR, our findings in Tables I indicate that significantly smaller models can also perform competitively on sequential path planning tasks. Moreover, SmallPlan’s SLMs are capable of optimizing side factors, as demonstrated by superior SPL and Dist. traveled metrics compared to GPT-4o. Note that, although some pretrained SLMs produce shorter Dist., their SR is significantly lower due to command execution failures that prematurely reduce Dist., making this outcome not optimal. Some pretrained SLMs occasionally fail to follow the required output format in our prompts. For example, the pretrained Qwen-2.5 sometimes produces responses in Chinese rather than English. This issue is resolved after training with SmallPlan.
We ultimately show that fine-tuned SLMs perform reliably even when the starting poses, target objects, and object distributions are altered between runs. We also attribute the strong performance of Phi-4-mini to its pretraining on complex reasoning tasks [53]. Llama-3, despite not having the highest SR, showing the largest SR improvement over its pretrained version (Table II), suggesting that this model can adapt and learn quite effectively. Qwen-2.5, while not matching the performance of ChatGPT or Phi4-mini, still show notable improvements over their pretrained versions.
TABLE II: SmallPlan’s SLMs on different training strategy.
Model Strategy SR % $\uparrow$ SPL % $\uparrow$ Dist. $\downarrow$ SRI % $\uparrow$
Qwen-2.5 SFT only 85.14 58.15 21.42 38.82
RL-SFT 83.43 48.07 19.87 36.03
RL-SFT ++ 78.86 59.31 11.49 28.58
Llama-3 SFT only 88.57 64.36 13.54 64.99
RL-SFT 93.14 66.10 17.65 73.50
RL-SFT ++ 74.30 57.83 9.69 38.41
Phi4-mini SFT only 91.43 58.40 18.26 29.02
RL-SFT 96.57 70.41 16.13 36.28
RL-SFT ++ 85.14 60.33 13.52 20.15
    “RL-SFT”=balanced reward, “RL-SFT ++”=strong distance penalty. The underline highlights the shortest Dist. within same SLM architecture. SRI denotes % SR increase compare to its pretrained version in Table I.
V-B2 Trade-offs and Side Factor Control with RL
Table II presents the performance of SLMs under different training strategies and reward penalties. We observe that applying RL does not always lead to higher SR. This is reasonable, as SFT alone is sufficient to transfer the reasoning capability from LLMs to SLMs effectively.
@@ V-B2 Trade-offs and Side Factor Control with RL
Notably, we highlight that the candidate SLM adapt to different reward penalties from training. For RL-SFT +, we set the $\lambda_{\text{efficiency}}$ twice to that of RL-SFT (0.6 vs. 0.3). This change result in a significant reduction in Dist. traveled by the SLMs while retain better SR compared to their pretrained versions, as shown in Table II. There exist trade-off between SR and Dist., as we observe that the SLMs tend to prematurely terminate tasks by calling done() instead of continuing to explore the environment. However, our approach paves the way for adaptation to different constraints and settings. While SFT trains the SLMs to reason by following the superior LLMs, it does not consider deployment constraints such as energy consumption. The SLMs trained with only SFT tend to re-prompt until reaching the maximum number of trials, whereas RL in our design allows early termination, reducing unnecessary computation. In real-time robotic applications on edge devices, each prompt incurs significant energy costs and action execution overhead, making efficient operation a key challenge. Thus, we consider a reasonable degree of the trade-off between SR. and Dist. acceptable. We conclude that RL plays a crucial role in controlling side factors and enables more resource-efficient planning for deployment.
V-B3 Few-shot SFT for Format Refinement.
@@
In Algorithm 1, we mentioned that SFT is applied before the main fine-tuning process for format refinement. In our experiment, removing this step results in unstable training dynamics and incoherent responses, as the scalar reward from RL alone is insufficient to guide learning. Therefore, we find that training from the few-shot fine-tuned models is far more effective than from the pretrained ones, as the SFT stage establishes a strong prior for LoRA adaptation and output format adherence.
Figure 4: Inference response for task ”find a microwave.” The system and user blocks show the task-level and step-level prompts, assistant shows the response by SmallPlan’s Phi-4-mini, env shows environment feedback. The bold texts indicate the information feed to the prompt, including the task and info from the scene graph and Voronoi graph on the right side.
V-B4 Out-of-domain Analysis
@@
Fig. 3 evaluates the logical reasoning capabilities of the SLMs on out-of-domain tasks beyond path planning. These tasks involve different types of sequential decision-making, allowing us to assess general reasoning ability. Our trained SLMs perform on par with, and sometimes exceed, their pretrained versions, as reflected by the consistent SR with only minor differences. This suggests that distillation from a superior LLM can help retain the general reasoning ability of smaller models, even across diverse domains. Given the complexity of cross-domain adaptation, we leave further investigation to future research works. This experiment merely highlights that our fine-tuned SLMs maintain strong generalization and are not overfitting to path planning tasks. The goal here is not to compare different SLMs against each other on their success rates, but to compare each trained SLM with its own pretrained version to demonstrate that our SLMs retain their general reasoning capability.
V-B5 Hyper-parameter Selection and Prompting
We try various hyper-parameter configurations during training, including RL rewards and the number of training epochs. Based on our experience, selecting these settings is largely a trial-and-error process, as no standardized method can guarantee optimal results. By default, we use $r_{\text{success}}=5$, $\lambda_{\text{success}}=0.3$, $\lambda_{\text{explore}}=0.1$, $\lambda_{\text{efficiency}}=0.3$, and $\lambda_{\text{format}}=0.1$. For penalization of Dist. traveled, we increase the efficiency weight to $\lambda_{\text{efficiency}}=0.6$. The learning rate is set at $1\times 10^{-5}$ for all training. Meanwhile, the number of epochs depends heavily on the SLM architecture, as different models vary in learning capacity and convergence. We use 4 epochs for Phi-4-mini and Llama-3, and 6 epochs for Qwen-2.5. Simply increasing or decreasing the number of epochs does not consistently improve performance. We tested multiple settings and selected the best-performers. Our models achieved their best performance with 3 to 6 epochs. Further tuning efforts can help to find optimal hyper-parameters and further improve the SLMs’ performance.
Fig. 4 captures the real-time response from SmallPlan’s Phi-4-mini for a specific task. In our prompt, we require the SLMs to provide the reason before proposing an action. This helps to improve the performance of the model. We observe that the SLM effectively follows the output format, as well as provide a concrete analysis and reasoning for its action from the provided context.
VI Limitations and Discussions
How small is enough? In addition to our proposed SLMs, which are around 3-4B params, we also tested smaller models, including DeepSeek-R1 version 1.5B and Llama-3 version 1B. However, DeepSeek-R1 exhibited the same issue observed in Qwen-2.5 (Section V-B1) and producing poor results. Llama-3 (1B) achieved a SR of 14.28% for pretrained and 51.42% for SmallPlan RL-SFT, increasing the SR by 3.6 times. While trainable and more efficient, we believe these results are still insufficient for practical use. However, with the rapid effort made by NVIDIA [54] and other researches [38], we are convinced that SLMs with even smaller params can soon be on par with larger models, given a proper training strategy.
@@ VI Limitations and Discussions
SFT and Distillation with Ground Truth Labels. Standard distillation typically combines the teacher outputs with supervision from true labels, enabling the student model to inherit both the teacher’s knowledge and align with task objectives. However, our Cross Entropy Loss for SFT relies solely on the teacher outputs. This is because our task - path planning in a simulation environment - provides scalar reward signals instead of differentiable ground-truth labels and cannot be directly used in standard loss functions for backpropagation. In our approach, this issue is resolved by leveraging not only SFT but RL, which takes simulation feedback alongside the teacher LLM’s outputs to optimize the candidate SLM.
@@
Choice of RL Algorithm. Recent works introduce different alternatives to PPO. For instance, DPO [55] eliminates the need for an explicit teacher model, making the training process more efficient while often achieving performance comparable to PPO. However, the DPO’s objective is fixed by pair-wise preferences and a KL-regularization term, making it difficult to integrate the additional reward components described in Section IV-B. In contrast, PPO accepts any scalar reward; we can mix multiple objectives, tune their weights, and monitor their individual effects during training. These reasons make PPO more feasible for our approach.
VII Conclusion
We introduce SmallPlan, a novel LLM-guided distillation framework that leverages SFT and RL to train SLMs for high-level path planning. We discuss our findings on various training strategies and show that the fine-tuned SLMs reach competitive performance compared to superior LLM such as GPT-4o. SmallPlan paves the way for practical deployment of language-driven planning systems on edge devices in real-world robotics.